OpenStudy (anonymous):

Find the vector equation for the tangent line to the curve of intersection of x^2+y^2=25 and y^2+z^2=20 at (3,4,2)

OpenStudy (anonymous):

I would take the gradient of each and go from there, but I would have to take the cross product with multiple variables which could take awhile. Any other way to do this?

OpenStudy (anonymous):

I think the intersections are just hyperbolas in the xz plane

OpenStudy (anonymous):

y^2 = 25-x^2 (25 -x^2) +z^2 = 20 z^2 -x^2 =-5

OpenStudy (anonymous):

That would be the intersection?

OpenStudy (anonymous):

sec...

OpenStudy (anonymous):

I started reading about this to help, and the way it appears to be done is exactly what you mentioned earlier -- taking the cross product of the gradients.

OpenStudy (anonymous):

I got <5+32t, -24t, sqrt(20)+48t> as the intersection, but I'm not sure where to go from there.

OpenStudy (anonymous):

those points... how'd you get them?

OpenStudy (anonymous):

I think you should be using (3,4,2)

OpenStudy (anonymous):

Instead of doing the cross product in variables I plugged in (3,4,2)... Found the cross product: <32,-24,48> and found the position vector: <5,0,sqrt(20)> Then just used the r0+t*r1 equation

OpenStudy (anonymous):

your slope is right, but not sure why you're not using the point of tangency for a point on your line...

OpenStudy (anonymous):

Ok I'm lost

OpenStudy (anonymous):

@Algebraic! the tangent vector passes through the point.

OpenStudy (anonymous):

Can you walk me through it? I know how to find the tangent vector once I have the equation for the intersection, but how do I get that?

OpenStudy (anonymous):

Here's what I have so far...\[ \ \ \ U(x,y,z) = x^2+y^2=25\\ \ \ \ V(x,y,z) = y^2+z^2=20\\ \text{Determine }\nabla U, \nabla V.\\ \ \ \ \nabla U(x,y,z)=2x\hat i+2y\hat j\\ \ \ \ \nabla V(x,y,z)=2y\hat j+2z\hat k\\ \text{Compute the gradients at }(3,4,2).\\ \ \ \ \nabla U(3,4,2)=6\hat i+8\hat j\\ \ \ \ \nabla V(3,4,2)=8\hat j+4\hat k\\ \text{The tangent }W\text{ is orthogonal to both }\nabla U, \nabla V\text{ at said point.}\\ \ \ \ W=\nabla U\times\nabla V =32\hat i-24\hat j+48\hat k \]

OpenStudy (anonymous):

That's what I got too! But, I'm not sure how to get the tangent vector equation from a parametric equation

OpenStudy (anonymous):

Then to determine the actual tangent vector equation you do...\[ \ \ \ T(t)=r_0+tr_1\\ \text{where }r_0=3\hat i+4\hat j+2\hat k\text{ and }r_1=W=32\hat i-24\hat j+48\hat k\\ \ \ \ T(t)=3\hat i+4\hat j+2\hat k+32t\hat i-24t\hat j+48t\hat k\\ \ \ \ \ \ \ \ \ \ =[3+32]\hat i+[4-24t]\hat j+[2+48t]\hat k\\ \text{... I think.} \]

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

3+32t rather

OpenStudy (anonymous):

Apologies again if its wrong and I know for a fact some of my terminology is off... I have relatively little experience with stuff past real analysis of a single variable and I'm just now in a Calculus I course.

OpenStudy (anonymous):

No worries!